# GraspDiffusion: Synthesizing Realistic Whole-body Hand-Object Interaction

### v1

@@ 3 GraspDiffusion
Table 1: Quantitative comparison on full-bodied generation We report the scores of current baselines on multiple evaluation metrics. FID [22] and KID [5] are used to measure the image quality, while CLIPScore [21] evaluates the alignment between the generated image and prompt.
@@
Table 2: Quantitative comparison on hand-object generation We report the results on both image quality metrics and successful hand-object grasps, showing that our method can outperform previous approaches.
@@ 4.2 Implementation Details
For the full-body grasping pipeline, we first train the body-pose diffusion model on the GRAB [65] dataset, which contains full-bodied humans interacting with various objects. We collect all frames that has more than 40 contacting vertices between the object and the subject’s right hand, and adopt the Adam optimizer [31] with a learning rate of $5\times 10^{-4}$. We train the model with batch size of 2,048 for 50k steps, using 2 RTX 6000 GPUS. For the diffusion schedule, we adapt a cosine noise schedule with $T=1000$.
To implement the scene generation pipeline, we train the two modules with the aforementioned custom datasets. For the hand refinement modules, we also add subsets from the Dex-YCB dataset [10] and the RHD dataset [84] to enhance the quality and quantity of hand-object images. Employing the Stable Diffusion v1.5 [58] as a base model with parameters frozen, we train the conditional modules with a constant learning rate of $10^{-4}$, for 200 epochs on four A100 GPUS which costs approximately 28 hours. For inference, we used a linear multistep scheduler [27] with 30 inference steps using a classifier-free guidance [24] of 3.5. We also support inference using personalized Stable Diffusion models other than the Stable Diffusion v1.5 model used during training, and display results with different models in Fig. 10.
Figure 8: Qualitative results We compare generated human-object interaction images generated by different methods using the same input object. Note that except for the first column, all images were based on the same human pose and object location created from our grasping pipeline. Despite being capable of generating high-quality images, other methods display erroneous interactions (e.g. multiple objects, object appearance distorted, color blending), while our pipeline can correctly convey the intention of the human-object grasping pipeline.
@@ 4.3 Quantitative Results
Methods Contact ratio $\uparrow$ Pose Valid Error $\downarrow$
GOAL-GNet [64] 0.461 0.504
COOP [81] 0.841 0.239
Ours 0.909 0.111
Table 3: Grasping pose evaluation We report the results on hand-object contact ratio and body pose validity between our method and prior methods, which displays our method’s significance in generating grasping poses.
Hand-grasp generation quality We also assess the quality of hand-object centric images from the hand refinement pipeline, based on both image quality and plausible hand-object pose. Along with the image quality metrics, to measure instances of successful hand-object contact, we adopt the contact evaluation setup in Affordance Diffusion [76] and utilize a widely used hand-object detector [60] to measure the object’s contact status. We compare our results with three baseline models : (1) a depth-based ControlNet, (2) HandRefiner [46], and (3) Affordance Diffusion [76]. We evaluate on a subset of the DexYCB dataset [10], and report the results in Table 2, which displays that our method is capable of outperforming previous methods in synthesizing hand-object images with accurate contact.
@@
Table 4: Ablation studies on architecture choice We compare the FID scores of our pipeline with cases of missing conditional modules. Our main setup, which utilizes all three modules, outperforms other cases.
3D pose evaluation To evaluate the plausibility of generated grasping poses for different objects and positions, we construct a test set of unseen objects distributed far from the original range of the training dataset. We choose 10 novel 3D objects [10] and 10 human body shapes, and for each object-body pair, we position the object at 64 random 3D positions, relative to the human body’s pelvis joint position. Specifically, the x-coordinate (the horizontal position in our paper) ranges from -0.5m to 0.5m, the y-coordinate (the vertical position in our paper) ranges from -0.8m to 0.8m, and the z-coordinate (the direction where the human model is facing) ranges from 0.0m to 0.8m, with the pelvis joint position as its origin.
To evaluate the validity of grasping poses, we measure the ratio of successful contact between the generated hands and the objects. To validate the plausibility of the generated body poses, we utilize VPoser [52], a variational human pose prior model. We measure the L2 loss of vertex reconstruction from the body poses as a pose valid error, given that an implausible body pose will result in a higher pose valid error. We compare our approach with two prior methods that are trained on the GRAB dataset [65] and support generating full-body grasps for different object translations; GOAL [64] and COOP [81]. For GOAL, we only evaluate the grasping pose generation with optimization (GNet) and set the x-coordinate of the object translation to 0 due to the fact that GOAL does not work when the objects are out of distribution in the horizontal plane [66]. We present the results in Table  3. The results demonstrate that our model is capable of generating authentic grasping poses for objects in various positions.
@@
In Table  1, we also present the quantitative results for the setting that does not uses the attention injection scheme, which alleviates the risk of generating erroneous human-object interactions. While the setting without attention injection has a slightly better FID / KID score, our full setting shows a substantial increase in CLIPScore, signifying that the generated images successfully adheres to the given interaction context.
@@
(a) Object Image
(b) 3D Mesh
(c) Result image
Figure 9: Example synthesis results from an object image We display generation results from a single object image, which was in turn used to generate a 3D model through TripoSR [67] and subsequently used as input to our pipeline.
@@
Figure 10: Example synthesis results from different Stable Diffusion models We display generation results from our pipeline with the same object and body pose, but with different personalized Stable Diffusion models that were acquired from CivitAI [1] and Huggingface [2].
@@ 4.4 Applications
By utilizing a 3D full-body grasping model generated from a simple object input, GraspDiffusion can provide a practical solution for AI practitioners who intend to use generative AI for their artwork such as advertisements, illustrations, and comic books. To alleviate the requirement of a 3D object mesh model, our pipeline also supports using 3D reconstruction models [67, 72, 45] that can recover 3D mesh models from a single image input. We display examples of image-based generations in Figure 9 and Figure 8 (bottom row). Also, our pipeline can support various personalized image domains, including (but not limited to) realistic, anime, pixel art style, and more. In Figure 10 we present results from the same 3D grasping pose, using diverse personalized text-to-image models to support different art styles and backgrounds. Further examples of such use cases will be provided in the Supplementary Material.
@@ 5 Discussions and Limitations
While GraspDiffusion can produce humans with detailed finger articulation and accurate object interaction, several samples exhibit discrepancies between the body’s skin texture and the refined hand’s skin texture. We account this issue to the shortage of balanced, high-quality data samples during training, and opt to construct additional interaction samples to facilitate high-quality generation. Also, GraspDiffusion can’t currently synthesize interactions that involve both hands or interactions between humans, which are scenarios that are also highly desired by practitioners. In the future, we aspire to extend our pipeline toward scene-level generation that involves interaction between multiple humans and objects.

### v2

@@ 4.1 Dataset Construction
Table 2: Quantitative comparison on hand-object generation
@@
(a) Input Object
(b) SDXL
(c) + Pose-ControlNet
(d) + Object Inpainting and HandRefiner
(e) Ours (w/o Attention Injection)
(f) Ours
Figure 8: Qualitative results. We compare HOI images generated by different methods based on a input object (first column). Note that except for the second column, all images were based on the same human pose and object location created from our grasping pipeline. While other methods display erroneous interactions (e.g. multiple objects, object appearance distorted, physically implausible interactions, color blending), which are marked with red segments, our scene-generation pipeline can correctly convey the interaction intention from the full-body grasping pipeline.
@@ 4.2 Implementation Details
For the full-body grasping pipeline, we train the body-pose diffusion model on the GRAB [68] dataset, which captures the full-bodied 3D SMPL-X interaction sequences with various objects. After downsampling the motion sequences of GRAB to 30 fps, we collect all frames that has more than 40 contacting vertices between the object and the subject’s right hand, to collect feasible grasps for the training. We then adopt the Adam optimizer [35] with a learning rate of $5\times 10^{-4}$. We train the model with batch size of 2,048 for 50k steps, using 2 RTX 6000 GPUS. For the diffusion schedule, we adapt a cosine noise schedule with $T=1000$.
To implement the scene generation pipeline, we train the two modules with the aforementioned custom datasets. Employing the Stable Diffusion v1.5 [62] as a base model with parameters frozen, we train the conditional modules with a constant learning rate of $10^{-4}$, for 200 epochs on four A100 GPUS which costs approximately 28 hours. For inference, we used a linear multistep scheduler [31] with 30 inference steps using a classifier-free guidance [26] of 3.5. We also support inference using personalized Stable Diffusion models other than the Stable Diffusion v1.5 model used during training, and display results with different models in Fig. 9.
@@ 4.3 Quantitative Results
Table 3: Grasping pose evaluation
Methods Contact ratio $\uparrow$ Pose Valid Error $\downarrow$ Displacement $\downarrow$
GOAL [69] 0.461 0.504 6.135
FLEX [70] 0.540 0.252 7.794
COOP [87] 0.841 0.239 4.679
Ours 0.909 0.111 2.696
Hand-grasp generation quality We also assess the quality of hand-centric images from the hand refinement pipeline, based on both image quality and plausible hand-object pose. To measure instances of successful hand-object contact, we adopt the contact evaluation setup in Affordance Diffusion [82] and utilize a widely used hand-object detector [64] to measure the object’s contact status. We compare our results with three baseline models : (1) a depth-based ControlNet, (2) HandRefiner [50], and (3) Affordance Diffusion [82]. We evaluated on a subset of the DexYCB dataset [11]. We report the results in Table 2, which shows that our method is capable of outperforming previous methods in creating hand-object images with accurate contact.
3D pose evaluation To evaluate the plausibility of grasping poses for different objects and positions, we constructed a test set of unseen objects distributed far from the original range of the training dataset. We choose 10 novel 3D objects from Dex-YCB [11] and 10 human body shapes, and for each pair, we position the object at 64 random 3D positions, relative to the human body’s pelvis joint position. Specifically, the x-coordinate (the horizontal position in our paper) ranges from -0.5m to 0.5m, the y-coordinate (the vertical position in our paper) ranges from -0.8m to 0.8m, and the z-coordinate (the direction where the human model is facing) ranges from 0.0m to 0.8m, with the pelvis joint position as its origin.
For pose evaluation, we utilize VPoser [56] to measure the L2 loss of vertex reconstruction from the body poses as a pose-valid error, given that an implausible body pose will result in a higher pose-valid error. We compare our approach with two prior methods that are trained on the GRAB dataset [68] and support generating full-body grasps for different object translations; GOAL [69] and COOP [87]. For GOAL, we only evaluate the grasping pose generation with optimization (GNet) and set the x-coordinate of the object translation to 0 due to the fact that GOAL does not work when the objects are out of distribution in the horizontal plane [70]. We present the results in Table  3. The results demonstrate that our model is capable of generating authentic grasping poses for objects in various positions.
Table 4: Ablation studies on architecture choice
@@
In Table 1, we also display results for the main setting without attention injection, which alleviates the risk of generating erroneous interactions. While the setting without attention injection has a slightly better FID / KID score, our full setting shows a substantial increase in CLIPScore, signifying that the generated images successfully adhere to the given interaction context.
@@
Figure 9: Example results from different models. We display generation results from our pipeline with the same object and body pose, but with different personalized Stable Diffusion models that were acquired from CivitAI [1] and Huggingface [2].
@@ 4.4 Applications
By utilizing a 3D full-body grasping model generated from a simple object input, GraspDiffusion can provide a practical solution for AI practitioners who intend to use generative AI for their artwork such as advertisements, illustrations, and comic books. To alleviate the requirement of a 3D object mesh model, our pipeline also supports using 3D reconstruction models [71, 77, 49] that can recover 3D mesh models from a single image input. Moreover, our pipeline can support various personalized image domains, including (but not limited to) realistic, anime, pixel art style, and more. In Figure 9 we present results from the same 3D grasping pose, using diverse personalized text-to-image models to support different art styles and backgrounds. Further examples of such use cases will be provided in the Supplementary Material.
@@ 5 Discussions and Limitations
While GraspDiffusion can produce humans with detailed finger articulation and accurate object interaction, several samples exhibit discrepancies between the body’s skin texture and the refined hand’s skin texture. We account this issue to the shortage of balanced, high-quality data samples during training, and opt to construct additional interaction samples to facilitate high-quality generation. In the future, we aspire to extend our pipeline toward scene-level generation that involves interaction between multiple humans and objects, and to better define interactions in the form of user-controllable text prompts. We also look forward to synthesize zero-shot interaction motions by leveraging image-to-video diffusion models [6, 81, 22, 75].